A bicycle wheel of diameter 80 cm accelerates uniformly from rest to an angular speed of

50 rad/s in 20s.

(a) Find the angular acceleration of the wheel.

2

(b) If a small stone is stack between the grooves of wheel, find the tangential and radial

acceleration of a stone.

(c) During this time, how many revolutions has the wheel turned?

(d) If the wheel turns through 20 revolutions in coming to rest, at what rate was the angular

velocity changing?


1
Expert's answer
2021-05-25T10:09:55-0400

(a)

"\alpha =\dfrac{\omega-\omega_0}{t},""\alpha =\dfrac{50\ \dfrac{rad}{s}}{20\ s}=2.5\ \dfrac{rad}{s^2}."

(b) We can find the tangential acceleration of the stone as follows:


"a_t=r\alpha=0.8\ m\cdot2.5\ \dfrac{rad}{s^2}=2\ \dfrac{m}{s^2}."

We can find the radial acceleration of the stone as follows:


"a_r=\dfrac{v^2}{r}=\dfrac{(\omega r)^2}{r}=\omega^2r,""a_r=(50\ \dfrac{rad}{s})^2\cdot0.8\ m=2000\ \dfrac{rad}{s^2}."

(c) We can find the number of revolutions of the wheel during the time "t=20\ s" from the kinematic equation:


"\theta=\omega_0t+\dfrac{1}{2}\alpha t^2,""\theta=0+\dfrac{1}{2}\cdot2.5\ \dfrac{rad}{s^2}\cdot(20\ s)^2=500\ rad.""n=500\ rad\cdot\dfrac{1\ rev}{2\pi\ rad}=80\ rev."

d) Let's first find the length traveled by the wheel during 20 revolutions:


"L=2\pi nr=2\pi\cdot20\cdot0.8\ m=100\ m."

Then, we can find the angle covered during deceleration:


"\theta=\dfrac{L}{r}=\dfrac{100\ m}{0.8\ m}= 125\ rad."

Finally, we can find the deceleration of the wheel (or the rate of change of angular velocity) from the kinematic equation:


"\omega^2=\omega_0^2+2\alpha\theta,""\alpha=\dfrac{\omega^2-\omega_0^2}{2\theta},""\alpha=\dfrac{0-(50\ \dfrac{rev}{s}\cdot\dfrac{2\pi\ rad}{1\ rev})^2}{2\cdot125\ rad}=-395\ \dfrac{rad}{s^2}."

The sign minus means that the wheel decelerates.